# Dual Magnetic and Electric Dipole Symmetry: Pseudo-Angular Momentum in Parity Space and the Electric Landé $g$-Factor

### v1

@@ Abstract
Electric dipole moments (EDMs) are sensitive probes of fundamental symmetries and central to searches for physics beyond the Standard Model. We present a symmetry-based description of EDMs, analogous to the Zeeman effect, that places magnetic and electric dipoles on equal footing under electromagnetic duality. In the hydrogenic problem, a static $\vec{B}$ (pseudovector) couples to the angular-momentum operator $\hat{\vec{J}}$ and reduces $SO(4)$ to the axial $SO(2)$ generated by $\hat{J}_{z}$. In contrast, a static $\vec{E}$ (polar vector) couples within a fixed $n$ manifold to a scaled Runge-Lenz operator $\hat{\vec{A}}_{\rm sc}$, mixes opposite parities, and preserves $SO(2)\times SO(2)$ generated by $\hat{J}_{z}$ (axial) and $\hat{A}_{{\rm sc},z}$ (dynamical). This motivates a pseudo-angular-momentum operator $\hat{\vec{J}}_{p}$, built from $\hat{\vec{A}}_{\rm sc}$, and an electric Landé factor $g_{E}$, so that the orbital dipole operator reads $\hat{\vec{d}}_{\rm orb}=g_{E}\,d_{B}\,\hat{\vec{J}}_{p}/\hbar$, with the natural electric unit $d_{B}=ea_{0}=\frac{2\mu_{B}}{c\alpha}$ ($a_{0}$ and $\mu_{B}$ are the Bohr radius and magneton respectively). As an example, Stark mixing of $2s$ and $2p_{m=0}$ in hydrogen yields $|\langle d_{\rm orb}\rangle|=3d_{B}$ (i.e. $g_{E}=3$). Following Ohanian’s magnetisation formalism for the Zeeman effect, we construct its electric dual; the wavefunction’s microscopic polarisation, which has a non-zero curl defining an effective magnetic probability current $\vec{J}_{m}=-\epsilon_{0}^{-1}\nabla\times\vec{P}$. The EDM expectation value becomes $\langle\hat{\vec{d}}_{tot}\rangle=-\frac{\epsilon_{0}}{2}\int\vec{r}\times\vec{J}_{m}\,d^{3}r=d_{B}\big[g_{E}\,\langle\hat{\vec{J}}_{p}\rangle/\hbar+g_{E}^{e}\,\langle\hat{\vec{S}}\rangle$$\hbar\big$], $g_{E}^{e}=\frac{2d_{\rm int}}{d_{B}}$. Here the spin operator, $\hat{\vec{S}}$, encodes any intrinsic EDM, $d_{\rm int}$, from symmetry-violating interactions, while $\hat{\vec{J}}_{p}$ captures the Stark-induced pseudo-angular momentum set by Runge-Lenz symmetry. This dual framework shows that induced EDMs arise from effective circulating probability magnetic currents, directly mirroring magnetic dipole moments arising from circulating probability electric currents.
@@ II Magnetic Dipole Moments
II.1 The Semi-Classical Bohr Magnetron
@@ III Electric Dipole Moments
So far, no particle has been observed to have a nonzero intrinsic EDM. In the Standard Model, such EDMs are predicted to exist but at levels far below current experimental sensitivity (for example, the electron EDM is expected to be about $10^{-38}\,\mathrm{e\cdot cm}$, whereas the neutron is about $10^{-31}\,\mathrm{e\cdot cm}$). State-of-the-art experiments constrain the electron EDM to be smaller than $1.3\times 10^{-29}\,\mathrm{e\cdot cm}$ and the neutron EDM below $1.0\times 10^{-26}\,\mathrm{e\cdot cm}$, with similar null results for the muon and proton. These increasingly tight bounds are crucial because a confirmed intrinsic EDM would signal new CP-violating physics beyond the Standard Model, possibly helping to explain the matter-antimatter asymmetry of the Universe. By contrast, the electric dipole moments that appear in the linear Stark effect of hydrogen or polar molecules are not intrinsic but induced. For example, in hydrogen they arise from parity mixing of nearly degenerate orbitals, such as the $2s$ and $2p_{0}$ states. In this case, the dipole operator acts on a pseudo-angular momentum degree of freedom, defined in the two-level parity subspace, and the resulting EDM depends on the external field and the mixing angle.
@@
III.1 Semi-Classical Model of a Quanta of Permanent Polarization and dipole moment
@@
Figure 1: Semi-classical representations of the Bohr EDM, $\vec{d}_{B}$. (a) Ideal semi-classical unit cell of quantised constant permanent polarization, $\Delta\vec{P}=\frac{e}{\pi a_{e}^{2}}\hat{z}$. In this model, the total angular momentum, and net electrical and magnetic charge are zero. However, because the curl of the polarization is non-zero, an effective magnetic current appears on the radial surface, which can be modelled as a quanta of a south and north pole ($\mp q_{m}=\mp\frac{h}{e}$) rotating in opposite directions, and is related to pseudo-spin (or pseudo-angular momentum) in parity space rather than position space. (b) Classical Bohr EDM, represented by two point charges $\pm e$ separated by a distance of the Bohr radius. (c) Schematic representing a semi-classical intrinsic Bohr EDM, connected to the position space spin vector, $\vec{S}$, related to an effective magnetic current of $\vec{I}_{m}$ at a radial distance of $a_{0}$.
Similar to the semi-classical atomic Bohr magneton calculation one can use a similar technique to derive the semi-classical atomic dual electrical analogue. To do this we conceive a toy model model of two quanta of opposite signed magnetic charge $\pm\frac{h}{e}$ as shown in Fig.1, rotating at the the electron Bohr radius, $a_{0}=\frac{\hbar}{m_{\mathrm{e}}c\alpha}$, creating a EDM. Here $\alpha=\frac{e^{2}}{4\pi\varepsilon_{0}\hbar c}$ is the fine structure constant, $\epsilon_{0}$ the permittivity of free space, $c$ the speed of light.
@@
In this model, the total net electrical and magnetic charge as well as angular momentum is zero. However a finite effective magnetic current exits, which we pose is not generated by mechanical orbital motion, but by pseudospin parity mixing. Given that $q_{m}=\frac{h}{e}$, the magnetic current at the boundary is given by,
$\vec{I}_{m}=-\frac{2q_{m}v}{2\pi a_{0}}\hat{\phi}=-\frac{e}{2\pi a_{0}^{2}m^{*}\epsilon_{0}\alpha c}(a_{0}m^{*}v)\hat{\phi}$ (14)
Here $m^{*}$ is the effective mass of the atomic particle related to the permanent dipole, which can vary from $0.01m_{e}$ to very large.
Assuming the total mass per unit cell (Fig.1(c)) is equivalent to the effective mass of the electron, then we assume $m^{*}\sim\frac{m_{e}}{2}$ (two magnetic charged particles), and equating $a_{0}m^{*}v\sim\hbar$, as in the derivation of the Bohr magneton in (5), then the magnetic current becomes, $\vec{I}_{m}=-\frac{e\hbar}{\pi a_{0}^{2}m_{e}\epsilon_{0}\alpha c}\hat{\phi}=-\frac{2\mu_{\mathrm{B}}}{\pi a_{0}^{2}\epsilon_{0}\alpha c}\hat{\phi}$. Thus, we can define the “Bohr EDM” in terms of the effective azimuthal magnetic current, via a left hand rule,
@@
the dual of the Bohr magneton in (5). This dipole moment is equivalent to two charges $\pm e$ separated by a distance of $a_{0}$ as shown in Fig. 1 (b), which also means $\vec{a}=a_{0}\hat{z}$, so Fig. 1(a) and 1(b) are equivalent.
@@
where $\vec{a}$ is any lattice vector and $V$ the unit-cell volume. This follows from an electron in a periodic potential, which gives the $U(1)$ gauge freedom (Bloch’s theorem); $\psi_{n\vec{k}}\!\to\!e^{-i\vec{k}\!\cdot\!\vec{a}}\psi_{n\vec{k}}$, which shifts the Berry phase by $2\pi$ and $\vec{P}$ by $e\vec{a}/V$. Here we consider a quasi-1D “cell” as shown in Fig.1 (c) of length $a$ (along $\hat{\vec{z}}$) and cross-section $A_{\!\mathrm{eff}}$, so the effective volume is, $V_{\!\mathrm{eff}}\;=\;A_{\!\mathrm{eff}}\,a$. In our construction $2\pi$ winding of the internal phase is equivalent to choosing $\vec{a}=a\hat{\vec{z}}$. Then from (16)
@@
This is equivalent to a single cell of a super cell in modern polarization theory [10], with a quanta of permanent polarization, $\Delta P_{z}$ of size on the order of the Bohr radius, where $A_{\!\mathrm{eff}}=\pi a_{0}^{2}$, as shown in Fig.1. Asserting that $\vec{a}=a_{0}\hat{z}$, the dipole moment can then be calculated from $\vec{d}_{B}=\int\Delta P_{z}d^{3}r=ea_{0}\hat{z}=d_{B}\hat{z}$, consistent with (15).
@@ III.2.1 Bohr Electric Dipole Moment from Spin
$=\frac{2d_{B}}{\hbar}\psi^{\dagger}(\vec{r})\hat{\vec{S}}\psi(\vec{r}),\qquad\hat{\vec{S}}\equiv\frac{\hbar}{2}\bm{\sigma}.$ (19)
where, $\bm{\sigma}$ denotes the vector of Pauli operators acting on the two-level parity space, $\vec{\Sigma}$ is the Pauli spin matrix vector and $\psi$ the Dirac or Pauli spinor. Adopting the dual (Weber) convention for magnetic current, the curl of $\vec{P}$ defines an effective probability magnetic current density,
@@ III.2.2 Intrinsic and Orbital Polarization
$\vec{P}_{\rm e}(\vec{r})=d_{int}\,\psi^{\dagger}(\vec{r})\,\bm{\Sigma}\,\psi(\vec{r})\;\;\approx\;\;\frac{2d_{int}}{\hbar}\,\vec{s}(\vec{r}),$ (24)
@@
where $\vec{s}(\vec{r})=\psi^{\dagger}(\vec{r})\frac{\hbar}{2}\bm{\sigma}\psi(\vec{r})$ is the nonrelativistic spin density and $g_{E}^{e}$ is the effective electric Landé g-factor of the electron. To understand how the electron behaves in an electric field we consider the interaction term in (1). To interpret this interaction physically, we examine its nonrelativistic limit, and begin by noting the electromagnetic field strength tensor:
@@
Since no electron EDM has been observed, it either vanishes or the intrinsic value $d_{\rm int}$ lies below current experimental sensitivity. Throughout this work we set $d_{\rm int}\simeq 0$, so the polarisation obtained from (31) reduces to $\vec{P}(\vec{r})\simeq\rho_{\rm charge}(\vec{r})\,\vec{r}$. To calculate orbital EDMs (e.g., the linear Stark effect), we use the effective magnetic probability current density, eq. (20), together with the polarisation equations (21) and (23). In the following we show that this relates to an effective angular-momentum like degree of freedom (a pseudo-angular momentum) that emerges when a system has an internal two- or few-level structure organised by a hidden symmetry.
@@ IV.2 Dual Ohanian Model for the $n=2$ Orbital
Applying the dual Ohanian model introduced in section III.2, we calculate the effective electron charge densities from the wavefunctions,

### v2

@@ Abstract
Electric dipole moments (EDMs) are sensitive probes of fundamental symmetries and central to searches for physics beyond the Standard Model. We present a symmetry-based description of EDMs, analogous to the Zeeman effect, that places magnetic and electric dipoles on equal footing under electromagnetic duality. In the hydrogenic problem, a static $\vec{B}$ (pseudovector) couples to the angular-momentum operator $\hat{\vec{J}}$ and reduces $SO(4)$ to the axial $SO(2)$ generated by $\hat{J}_{z}$. In contrast, a static $\vec{E}$ (polar vector) couples within a fixed $n$ manifold to a scaled Runge-Lenz operator $\hat{\vec{A}}_{\rm sc}$, mixes opposite parities, and preserves $SO(2)\times SO(2)$ generated by $\hat{J}_{z}$ (axial) and $\hat{A}_{{\rm sc},z}$ (dynamical). This motivates a pseudo-angular-momentum operator $\hat{\vec{J}}_{p}$, built from $\hat{\vec{A}}_{\rm sc}$, and an electric Landé factor $g_{E}$, so that the orbital dipole operator reads $\hat{\vec{d}}_{\rm orb}=g_{E}\,d_{B}\,\hat{\vec{J}}_{p}/\hbar$, with the natural electric unit $d_{B}=ea_{0}=\frac{2\mu_{B}}{c\alpha}$ ($a_{0}$ and $\mu_{B}$ are the Bohr radius and magneton respectively). As an example, Stark mixing of $2s$ and $2p_{m=0}$ in hydrogen yields $|\langle d_{\rm orb}\rangle|=3d_{B}$ (i.e. $g_{E}=3$). Following Ohanian’s magnetisation formalism for the Zeeman effect, we construct its electric dual; the wavefunction’s microscopic polarisation, which has a non-zero curl defining an effective magnetic probability current $\vec{J}_{m}=-\epsilon_{0}^{-1}\nabla\times\vec{P}$. The EDM expectation value becomes $\langle\hat{\vec{d}}_{tot}\rangle=-\frac{\epsilon_{0}}{2}\int\vec{r}\times\vec{J}_{m}\,d^{3}r=d_{B}\big[g_{E}\,\langle\hat{\vec{J}}_{p}\rangle/\hbar+g_{E}^{e}\,\langle\hat{\vec{S}}\rangle$$\hbar\big$], $g_{E}^{e}=\frac{2d_{\rm int}}{d_{B}}$. Here the spin operator, $\hat{\vec{S}}$, encodes any intrinsic EDM, $d_{\rm int}$, from symmetry-violating interactions, while $\hat{\vec{J}}_{p}$ captures the Stark-induced pseudo-angular momentum set by Runge-Lenz symmetry. In this dual framework, an induced EDM, which physically originates from charge displacement via parity mixing, can be equivalently represented by an effective circulating probability magnetic current, mirroring the probability electric-current representation of magnetic dipole moments.
@@ II Magnetic Dipole Moments
II.1 The Semi-Classical Bohr Magneton
@@ III Electric Dipole Moments
So far, no particle has been observed to have a nonzero intrinsic EDM. In the Standard Model, such EDMs are predicted to exist but at levels far below current experimental sensitivity (for example, the electron EDM is expected to be about $10^{-38}\,\mathrm{e\cdot cm}$, whereas the neutron is about $10^{-31}\,\mathrm{e\cdot cm}$). State-of-the-art experiments constrain the electron EDM to be smaller than $4.1\times 10^{-30}\,\mathrm{e\cdot cm}$ [18] and the neutron EDM below $1.8\times 10^{-26}\,\mathrm{e\cdot cm}$ [19], with similar null results for the muon and proton. These increasingly tight bounds are crucial because a confirmed intrinsic EDM would signal new CP-violating physics beyond the Standard Model, possibly helping to explain the matter-antimatter asymmetry of the Universe. By contrast, the electric dipole moments that appear in the linear Stark effect of hydrogen or polar molecules are not intrinsic but induced. For example, in hydrogen they arise from parity mixing of nearly degenerate orbitals, such as the $2s$ and $2p_{0}$ states. In this case, the dipole operator acts on a pseudo-angular momentum degree of freedom, defined in the two-level parity subspace, and the resulting EDM depends on the external field and the mixing angle.
@@
III.1 Semi-Classical Mnemonic for a Quantum of Permanent Polarization and Dipole Moment
@@
Figure 1: Semi-classical representations of the Bohr EDM scale, $\vec{d}_{B}$. (a) Idealized quasi-1D “unit cell” of quantized permanent polarization, $\Delta\vec{P}=\frac{e}{\pi a_{0}^{2}}\hat{z}$, for which $\nabla\times\vec{P}\neq 0$ implies an effective magnetic current on the radial surface (a dual representation of the polarization texture, not a literal magnetic-charge flow). For mnemonic purposes this surface current can be pictured as two fictitious opposite magnetic charges $\mp q_{m}=\mp\frac{h}{e}$ circulating in opposite senses, associated with pseudo-spin (pseudo-angular momentum) in parity space rather than mechanical orbital motion. (b) Classical Bohr EDM: two point charges $\pm e$ separated by the Bohr radius $a_{0}$. (c) Schematic of an intrinsic Bohr-scale EDM connected to the position-space spin vector $\vec{S}$, represented equivalently by an effective azimuthal magnetic current $\vec{I}_{m}$ at radius $\sim a_{0}$.
We define the Bohr EDM scale as
$d_{B}\;\equiv\;ea_{0},$ (14)
which is the natural atomic dipole unit. In the Coulomb problem this scale is consistent with the hydrogenic $SO(4)$ (Runge–Lenz) structure within a fixed $n$-manifold, and it also has an immediate classical interpretation: it is the dipole moment of two point charges $\pm e$ separated by a distance $a_{0}$ [Fig. 1(b)]. Furthermore, the same dipole unit appears in a quasi-1D “cell” formulation governed by modern polarization theory, when the size of the cell spacing is of order $a_{0}$.
The following toy model is included only as a heuristic mnemonic (not as a physical derivation) to connect this dipole scale to a dual-current picture. Similar to the semi-classical Bohr magneton estimate, we imagine an effective surface magnetic current associated with the curl of the polarization. For mnemonic purposes, this may be visualized as two fictitious opposite “magnetic charges” $\pm\frac{h}{e}$ circulating at the Bohr radius $a_{0}=\frac{\hbar}{m_{\mathrm{e}}c\alpha}$, where $\alpha=\frac{e^{2}}{4\pi\varepsilon_{0}\hbar c}$.
@@
In this picture, the net electric charge, net magnetic charge, and total angular momentum of the cell are zero, but an effective magnetic surface current exists, which we associate with pseudo-spin (parity mixing) rather than mechanical orbital motion. With $q_{m}=\frac{h}{e}$, the effective magnetic current at the boundary is written as
$\vec{I}_{m}=-\frac{2q_{m}v}{2\pi a_{0}}\hat{\phi}=-\frac{e}{2\pi a_{0}^{2}m^{*}\epsilon_{0}\alpha c}(a_{0}m^{*}v)\hat{\phi}$, (15)
where $m^{*}$ is an effective inertia parameter for the internal (parity-space) degree of freedom. For a Bohr-scale dimensional estimate we take the usual quantization scale $a_{0}m^{*}v\sim\hbar$ with $m^{*}\sim\mathcal{O}(m_{e})$ (the precise value of $m^{*}$ is not used elsewhere in the manuscript). This gives
$\vec{I}_{m}\approx-\frac{e\hbar}{\pi a_{0}^{2}m_{e}\epsilon_{0}\alpha c}\hat{\phi}=-\frac{2\mu_{\mathrm{B}}}{\pi a_{0}^{2}\epsilon_{0}\alpha c}\hat{\phi}$, (16)
and we may express the Bohr EDM in terms of the effective azimuthal magnetic current (left-hand rule) as
@@
the electromagnetic dual of the Bohr magneton in (5). Thus Fig. 1(a) and Fig. 1(b) represent the same Bohr-scale dipole unit: $\vec{a}=a_{0}\hat{z}$ implies $d_{B}=ea_{0}$.
@@
where $\vec{a}$ is any lattice vector and $V$ the unit-cell volume. This follows from an electron in a periodic potential, which gives the $U(1)$ gauge freedom (Bloch’s theorem); $\psi_{n\vec{k}}\!\to\!e^{-i\vec{k}\!\cdot\!\vec{a}}\psi_{n\vec{k}}$, which shifts the Berry phase by $2\pi$ and $\vec{P}$ by $e\vec{a}/V$. Here we consider a quasi-1D “cell” as shown in Fig. 1 of length $a$ (along $\hat{\vec{z}}$) and cross-section $A_{\!\mathrm{eff}}$, so the effective volume is $V_{\!\mathrm{eff}}=A_{\!\mathrm{eff}}a$. In our construction $2\pi$ winding of the internal phase is equivalent to choosing $\vec{a}=a\hat{\vec{z}}$. Then from (18)
@@
This is equivalent to a single cell of a supercell in modern polarization theory [10], with a quantum of permanent polarization $\Delta P_{z}$ set by $A_{\!\mathrm{eff}}=\pi a_{0}^{2}$, as shown in Fig. 1. With $\vec{a}=a_{0}\hat{z}$, the dipole moment follows as $\vec{d}_{B}=\int\Delta P_{z}\,d^{3}r=ea_{0}\hat{z}=d_{B}\hat{z}$, consistent with (17).
For completeness when discussing the hydrogenic problem, we note that, for a finite-mass nucleus, the mass parameter in the internal (relative-coordinate) Coulomb problem should be understood as the reduced mass
$\mu\;=\;\frac{m_{e}m_{N}}{m_{e}+m_{N}},$ (20)
rather than the electron mass $m_{e}$, so that the Bohr radius is replaced by
$a_{\mu}\;=\;\frac{4\pi\epsilon_{0}\hbar^{2}}{\mu Ze^{2}},$ (21)
and hence the Bohr dipole scale used throughout becomes $d_{B}=ea_{\mu}$ (rather than $ea_{0}$ defined using $m_{e}$). This substitution rescales lengths and energies but does not modify the underlying symmetry structure: the $SO(4)\simeq SU(2)\times SU(2)$ algebra and associated commutation relations remain unchanged (see Appendix C), with $m\rightarrow\mu$ in the Hamiltonian. In practice, for hydrogen the correction is very small,
$\frac{\mu}{m_{e}}\;=\;\frac{1}{1+m_{e}/m_{p}}\;\approx\;0.99946,$ (22)
so $a_{\mu}$ (and thus $d_{B}$) differs from the $m_{e}$ value by only $\sim 5.4\times 10^{-4}$ (about $0.054\%$), and either convention may be adopted depending on the desired level of precision.
@@ III.2.1 Bohr Electric Dipole Moment from Spin
$=\frac{2d_{B}}{\hbar}\psi^{\dagger}(\vec{r})\hat{\vec{S}}\psi(\vec{r}),\qquad\hat{\vec{S}}\equiv\frac{\hbar}{2}\vec{\sigma}.$ (24)
where, $\vec{\sigma}$ denotes the vector of Pauli operators acting on the two-level parity space, $\vec{\Sigma}$ is the Pauli spin matrix vector and $\psi$ the Dirac or Pauli spinor. Adopting the dual (Weber) convention for magnetic current, the curl of $\vec{P}$ defines an effective probability magnetic current density,
@@ III.2.2 Intrinsic and Orbital Polarization
$\vec{P}_{\rm e}(\vec{r})=d_{int}\,\psi^{\dagger}(\vec{r})\,\vec{\Sigma}\,\psi(\vec{r})\;\;\approx\;\;\frac{2d_{int}}{\hbar}\,\vec{s}(\vec{r}),$ (29)
@@
where $\vec{s}(\vec{r})=\psi^{\dagger}(\vec{r})\frac{\hbar}{2}\vec{\sigma}\psi(\vec{r})$ is the nonrelativistic spin density and $g_{E}^{e}$ is the effective electric Landé g-factor of the electron. To understand how the electron behaves in an electric field we consider the interaction term in (1). To interpret this interaction physically, we examine its nonrelativistic limit, and begin by noting the electromagnetic field strength tensor:
@@
Since no electron EDM has been observed, it either vanishes or the intrinsic value $d_{\rm int}$ lies below current experimental sensitivity. Throughout this work we set $d_{\rm int}\simeq 0$, so that the polarization obtained from (36) reduces to $\vec{P}(\vec{r})\simeq\rho_{\rm charge}(\vec{r})\,\vec{r}$. To calculate orbital EDMs (e.g., the linear Stark effect), we use the effective magnetic probability current density, Eq. (25), together with the polarization relations (26) and (28). Importantly, the induced orbital EDM is produced physically by parity mixing (charge displacement); the associated magnetic-current description is a topological field equivalence (a dual representation of the polarization texture), not an additional causal mechanism. We further show that this charge displacement can be organized in terms of an effective angular-momentum like degree of freedom, a pseudo-angular momentum in parity space, that emerges when the system possesses an internal two- (or few-) level structure governed by a hidden symmetry.
@@ IV.2 Dual Ohanian Model for the $n=2$ Orbital
Applying the dual Ohanian model introduced in section III.2, we calculate the effective electron charge densities from the wavefunctions. Note, here the electron coordinates are relative to the nucleus and that the associated microscopic “effective” fields are internal fields defined in the atomic rest frame,